Gorillaz1002:

Rick got 7 over 8 pound of honey from a hive. He calculated the quotient of 7 over 8 ÷ 5 over 6. Which statement shows the quotient and gives the most likely explanation of what Rick was trying to determine? He can make 35 over 42 of a bottle of honey with 5 over 6 pound of honey in each bottle. He can get 35 over 42 pound of honey if he divides the honey into 5 over 6-pound portions. He can make 1 and 2 over 40 bottles of honey with 5 over 6 pound of honey in each bottle. He can get 1 and 2 over 40 pounds of honey if he divides the honey into 5 over 6-pound portions. AND -NO- ¨BRAINLY¨

jhonyy9:

\[\left( \frac{ 7 }{ 8 } \right)/ \left( \frac{ 5 }{ 6 } \right) = \left( \frac{ 7 }{ 8 } \right)×\left( \frac{ 6 }{ 5 } \right)\] hope this help you

Gorillaz1002:

thx

jhonyy9:

np anytime yw